# LAENeRF: Local Appearance Editing for Neural Radiance Fields

### v1

@@ Abstract
Due to the omnipresence of Neural Radiance Fields (NeRFs), the interest towards editable implicit 3D representations has surged over the last years. However, editing implicit or hybrid representations as used for NeRFs is difficult due to the entanglement of appearance and geometry encoded in the model parameters. Despite these challenges, recent research has shown first promising steps towards photorealistic and non-photorealistic appearance edits. The main open issues of related work include limited interactivity, a lack of support for local edits and large memory requirements, rendering them less useful in practice. We address these limitations with LAENeRF, a unified framework for photorealistic and non-photorealistic appearance editing of NeRFs. To tackle local editing, we leverage a voxel grid as starting point for region selection. We learn a mapping from expected ray terminations to final output color, which can optionally be supervised by a style loss, resulting in a framework which can perform photorealistic and non-photorealistic appearance editing of selected regions. Relying on a single point per ray for our mapping, we limit memory requirements and enable fast optimization. To guarantee interactivity, we compose the output color using a set of learned, modifiable base colors, composed with additive layer mixing. Compared to concurrent work, LAENeRF enables recoloring and stylization while keeping processing time low. Furthermore, we demonstrate that our approach surpasses baseline methods both quantitatively and qualitatively.
@@ 1 Introduction
Novel view synthesis has been completely revolutionized by Neural Radiance Fields (NeRFs) [28]. NeRFs enable high-fidelity reconstruction of a 3D scene from a set of input images and their camera poses, building on differentiable volume rendering. Recent methods have successfully applied NeRFs to dynamic scenes [34, 56, 57], large-scale scene reconstruction [4, 21, 39] and varying lighting conditions [6, 26]. Local appearance editing of these learned 3D scene representations remains relatively underexplored. The implicit representation used in NeRFs in the form of a multi-layer perceptron (MLP) is the main challenge, causing non-local effects when a single parameter is modified, Distilled feature fields [19] and per-image 2D masks [23] have been suggested as mechanisms to facilitate local edits for NeRFs. However, both of these methods frequently introduce artefacts in the non-edited regions. Other editing approaches support recoloring a NeRF by remapping individual colors [11, 20, 49], try to extract modifiable material quantities for re-rendering [5, 50], or apply style transfer [13, 54]. Virtually all approaches in these domains do not support controllable local edits, i.e., they always also introduce global changes, which constrains their viability to the theoretical domain and impedes their applicability in practical contexts. At the same time, most approaches struggle with high memory requirements and long compute times, further hampering their use. Ultimately, there is currently no method that enables simultaneous style transfer and interactive recoloring.
To address the previously discussed limitations, we propose LAENeRF, a method for local appearance editing of pre-trained NeRFs. Choosing NeRFShop [16] and Instant-NGP (iNGP[29] as building blocks, we use a 3-dimensional grid, a subset of iNGP’s occupancy grid, as our primitive for selecting scene content. Due to the region growing procedure inherited from NeRFShop, which relies on a growing queue storing direct neighbors, we can model smooth transitions to content adjacent to our selection, resulting in more visually appealing edits. Inspired by previous recoloring approaches [11, 20], we introduce a novel NeRF-like module designed to learn a palette-based decomposition of colors within a selected region. In contrast to previous work, we estimate a per-ray termination point resulting in a point cloud which represents the editable region. This design decision reduces memory requirements and increases performance drastically. We feed these points into our neural LAENeRF module, which learns a palette-based decomposition by jointly optimizing two MLPs and a set of base colors to reconstruct the selected region (see Fig. 1). As LAENeRF learns a function in 3D space, we can implicitly ensure multi-view consistency and prune outliers. The learned set of colors may be modified after optimization to enable interactive recoloring.
@@ 3 Preliminaries
which can in turn be used to compute the estimated ray termination
@@
NeRFShop [16] leverages a voxel grid, akin to the occupancy grid $\bm{\mathcal{O}}$ of iNGP [29], to select arbitrary content in iNGP-based radiance fields. The occupancy grid of iNGP is an acceleration structure, discretizing the bounded domain into uniformly sized voxels, with a value of 0 or 1 depending on the density. We utilize this concept with our edit grid $\bm{\mathcal{E}}$. To facilitate region selection, the user scribbles on the projection of the 3D scene on the screen. Subsequently, for each selected pixel, a ray is cast and the estimated ray termination ${\bm{x}}_{\text{term}}$ is computed with Eq. 5. Next, we map each ${\bm{x}}_{\text{term}}$ to the nearest voxel in our edit grid $\bm{\mathcal{E}}$ and set the corresponding bit in the underlying bitfield. For intuitive selection, we support region growing from the selected voxels, by adding the neighboring voxels to a growing queue ${\mathcal{G}}$. During region growing, we add the current voxel to $\bm{\mathcal{E}}$ and its neighbors to ${\mathcal{G}}$ if the occupancy grid is set. Through this workflow, we offer an intuitive method for content selection within iNGP’s hybrid representation.
Figure 2: Overview of our method, LAENeRF. Given the estimated ray terminations ${\bm{x}}_{\text{term}}$ for a region specified by $\bm{\mathcal{E}}$, we learn a mapping from these inputs to barycentric weights ${\hat{{\bm{w}}}}$ and offsets ${\hat{\bm{\delta}}}$. We compose the per-ray color ${\hat{{\bm{c}}}}$ using these intermediate outputs and a learnable color palette ${\hat{\bm{\mathcal{P}}}}$ and supervise with a content loss ${\mathcal{L}_{\text{content}}}$ and optional style losses ${\mathcal{L}_{\text{style}}},{\mathcal{L}_{\text{TV}}},{\mathcal{L}_{\text{depth-disc}}}$ to obtain a unified approach which supports recoloring and stylization.
@@ 4 LAENeRF
Our key insight is that we can reduce memory and computational requirements by optimizing a small, NeRF-like network given the estimated ray termination ${\bm{x}}_{\text{term}}$. Given the implicit assumption that our pre-trained NeRF accurately reconstructs the scene geometry, our formulation is inherently view-consistent. We present our network architecture in Fig. 2. ${\bm{x}}_{\text{term}}$ is featurized using a trainable multi-resolution hash grid [29]. The encoded input $\xi({\bm{x}}_{\text{term}})$ is subsequently passed to shallow MLPs to predict per-ray barycentric weights ${\hat{{\bm{w}}}}\in[0,1]^{{N_{{\hat{\bm{\mathcal{P}}}}}}}$ and offsets ${\hat{\bm{\delta}}}\in[-1,1]^{3}$, where ${N_{{\hat{\bm{\mathcal{P}}}}}}$ denotes the size of the learnable color palette ${\hat{\bm{\mathcal{P}}}}\in\mathbb{R}^{{N_{{\hat{\bm{\mathcal{P}}}}}}\times 3}$. With our predicted intermediate values, we compose our output color as
@@
To allow the offsets to model view-dependent effects, we featurize ${\bm{d}}$ using a spherical harmonics encoding [8, 44] and use this as an additional input to our offset network.
@@
As can be seen in Fig. 2LAENeRF learns a mapping from ${\bm{x}}_{\text{term}}$ to estimated output color ${\hat{{\bm{c}}}}$. However, we only require ${\bm{x}}_{\text{term}}$ of rays ${\bm{r}}$ which intersect $\bm{\mathcal{E}}$. To correctly handle occlusions and mitigate errors due to the edit grid resolution, we require the alpha accumulated inside the edit grid to be larger than a certain threshold, ${\tau_{\text{edit}}}=0.5$. To obtain accurate depth estimations, requiring the full accumulation of alpha, we simultaneously perform raymarching through both $\bm{\mathcal{E}}$ and the occupancy grid $\bm{\mathcal{O}}$, computing $0pt$ within the occupancy grid.
@@
As is common in 2D image style transfer [9, 17], we use separate losses for content and style, in our case:
${\mathcal{L}_{\text{content}}}$ $=\|{\hat{{\bm{c}}}}-{\bm{c}}\|_{2}^{2},$ (7)
@@
where ${{\bm{s}}}$ denotes an arbitrary style image and ${\bm{G}}(\cdot)$ denotes the gram matrix [9] of a feature extracted from a semantic encoder, e.g. VGG19 [36]. Crucially, we can perform photorealistic recoloring by setting ${\lambda}_{\text{style}}=0$, where LAENeRF learns to reconstruct the selected region.
@@
We notice that when we perform non-photorealistic stylization of 3D regions using only ${\mathcal{L}_{\text{content}}}$ and ${\mathcal{L}_{\text{style}}}$, small structures are often eliminated in favor of a consistent stylization. In addition, as imperfect geometry reconstruction from our pre-trained $\Theta_{\text{NeRF}}$ leads to noise in ${\bm{x}}_{\text{term}}$, we need additional regularization to encourage smooth, low-noise outputs. To facilitate detailed, geometry-aware stylization, we introduce two novel losses, which condition LAENeRF on the estimated geometry of $\Theta_{\text{NeRF}}$. First, we encourage LAENeRF to limit noise in regions where there are no depth discontinuities leveraging a depth guidance image:
@@
where $\nabla{\hat{{\bm{c}}}}$ denotes the image gradients of ${\hat{{\bm{c}}}}$ in $x/y$-direction, i.e. $(\nabla{\hat{{\bm{c}}}})_{i,j}=\|{\hat{{\bm{c}}}}_{i+1,j}-{\hat{{\bm{c}}}}_{i,j}\|_{2}^{2}+\|{\hat{{\bm{c}}}}_{i,j+1}-{\hat{{\bm{c}}}}_{i,j}\|_{2}^{2}$. This loss term remedies noisy prediction for ${\bm{x}}_{\text{term}}$, but does not preserve fine, geometric structures to a sufficient extent. Hence, we introduce another loss, which is minimized when image gradients are placed in regions of depth discontinuities:
@@
To prevent extreme solutions with a high-frequency offset function dominating the color prediction [1], we regularize the offsets with
@@
To edit the apperance of a pre-trained radiance field, we first extract a training dataset for LAENeRF. After optimization, we can compose a modified training dataset for fine-tuning $\Theta_{\text{NeRF}}$. We perform blending of the target colors ${\mathcal{C}}({\bm{r}})$ and LAENeRF’s output ${\hat{{\bm{c}}}}$ using the per-ray accumulated alpha as the blending weights.
@@
where ${\tau_{\text{dist}}}$ is a hyperparameter controlling the size of the transition, which we usually set to $1\times 10^{-2}$, depending on the scale of the scene. As can be seen in Fig. 3, we interpolate the original color palette ${\hat{\bm{\mathcal{P}}}}$ and the modified color palette ${\hat{\bm{\mathcal{P}}}}_{\text{mod}}$ based on ${d_{\text{trans}}}$ to get a realistic color transition. For stylization, we additionally introduce a separate loss, which adds to the content loss when ${d_{\text{trans}}}\in(0,1]$:
@@
We build our approach on torch-ngp [40], a PyTorch variant of iNGP [29]. For our edit grid $\bm{\mathcal{E}}$, we use a resolution of $128^{3}$, the same as for the occupancy grid. We train LAENeRF for $1\times 10^{5}$ iterations with previews available after $\sim 20$s, and then distill to $\Theta_{\text{NeRF}}$ for $7\times 10^{4}$ iterations, which takes no longer than 5 minutes in total on an NVIDIA RTX 4090. For LAENeRF, we use a learning rate of $1\times 10^{-3}$ for all components, except for ${\hat{\bm{\mathcal{P}}}}$, where we use a learning rate of $1\times 10^{-2}$. We initialize ${N_{{\hat{\bm{\mathcal{P}}}}}}=8$, a number which is usually too high for any scenario. However, we note that when we want to obtain a smooth transition utilizing ${\mathcal{L}_{\text{smooth}}}$, the increased number of base colors is vital. $1.5\times 10^{3}$ iterations before training is finished, we remove color palettes which do not contribute significantly. For our style loss, we use a VGG19 backbone [36] and utilize features from conv5, conv6, conv7 for computing ${\mathcal{L}_{\text{style}}}$. When performing stylization, we use ${\lambda}_{\text{style}}=1.3\times 10^{2},{\lambda}_{\text{TV}}=1\times 10^{-4},{\lambda}_{\text{depth-disc}}=5\times 10^{-4},{\lambda}_{\text{weight}}=1\times 10^{-7},{\lambda_{\text{offset}}}=5\times 10^{-5},{\lambda}_{\text{smooth}}=1\times 10^{-3}$. If we want to perform photorealistic recoloring, we set ${\lambda}_{\text{style}}={\lambda}_{\text{TV}}={\lambda}_{\text{depth-disc}}=0$.

### v2

@@ Abstract
Due to the omnipresence of Neural Radiance Fields (NeRFs), the interest towards editable implicit 3D representations has surged over the last years. However, editing implicit or hybrid representations as used for NeRFs is difficult due to the entanglement of appearance and geometry encoded in the model parameters. Despite these challenges, recent research has shown first promising steps towards photorealistic and non-photorealistic appearance edits. The main open issues of related work include limited interactivity, a lack of support for local edits and large memory requirements, rendering them less useful in practice. We address these limitations with LAENeRF, a unified framework for photorealistic and non-photorealistic appearance editing of NeRFs. To tackle local editing, we leverage a voxel grid as starting point for region selection. We learn a mapping from expected ray terminations to final output color, which can optionally be supervised by a style loss, resulting in a framework which can perform photorealistic and non-photorealistic appearance editing of selected regions. Relying on a single point per ray for our mapping, we limit memory requirements and enable fast optimization. To guarantee interactivity, we compose the output color using a set of learned, modifiable base colors, composed with additive layer mixing.selection. Compared to concurrent work, LAENeRF enables recoloring and stylization while keeping processing time low. Furthermore, we demonstrate that our approach surpasses baseline methods both quantitatively and qualitatively.
@@ 1 Introduction
Novel view synthesis has been completely revolutionized by Neural Radiance Fields (NeRFs) [28]. NeRFs enable high-fidelity reconstruction of a 3D scene from a set of input images and their camera poses, building on differentiable volume rendering. Recent methods have successfully applied NeRFs to dynamic scenes [34, 55, 56], large-scale scene reconstruction [4, 21, 39] and varying lighting conditions [6, 26]. Local appearance editing of these learned 3D scene representations remains relatively underexplored. The implicit representation used in NeRFs in the form of a Multi-Layer Perceptron (MLP) is the main challenge, causing non-local effects when a single parameter is modified. Distilled feature fields [19] and per-image 2D masks [23] have been suggested to facilitate local edits for NeRFs. However, both of these methods frequently introduce artefacts in the non-edited regions. Other editing approaches support recoloring a NeRF by remapping individual colors [11, 20, 48], try to extract modifiable material quantities for re-rendering [5, 49], or apply style transfer [13, 53]. Virtually all approaches in these domains do not support controllable local edits, i.e., they always also introduce global changes, which constrains their viability to the theoretical domain and impedes their applicability in practice. At the same time, most approaches struggle with high memory requirements and long compute times, further hampering their use. Ultimately, no method currently enables simultaneous style transfer and interactive recoloring.
To address the previously discussed limitations, we propose LAENeRF, a method for local appearance editing of pre-trained NeRFs. Choosing NeRFShop [16] and Instant-NGP (iNGP[29] as building blocks, we use a 3-dimensional grid, a subset of iNGP’s occupancy grid, as our primitive for selecting scene content. Due to the region growing procedure inherited from NeRFShop, which relies on a growing queue storing direct neighbors, we can model smooth transitions to content adjacent to our selection, resulting in more visually appealing edits.
Inspired by previous recoloring approaches [11, 20], we introduce a novel NeRF-like module designed to learn a palette-based decomposition of colors within a selected region. In contrast to previous work, we estimate a per-ray termination point resulting in a point cloud which represents the editable region. This design decision reduces memory requirements and increases performance drastically. We feed these points into our neural LAENeRF module, which learns a palette-based decomposition by jointly optimizing two MLPs and a set of base colors to reconstruct the selected region (see Fig. 1). As LAENeRF learns a function in 3D space, we can implicitly ensure multi-view consistency and prune outliers. The learned set of colors may be modified after optimization to enable interactive recoloring.
@@ 3 Preliminaries
which we use to compute the estimated ray termination
@@
NeRFShop [16] leverages a voxel grid, akin to the occupancy grid $\bm{\mathcal{O}}$ of iNGP [29], to select arbitrary content in iNGP-based radiance fields. The occupancy grid of iNGP is an acceleration structure, discretizing the bounded domain into uniformly sized voxels, with a value of 0 or 1 describing the expected density within this voxel. We utilize this concept with our edit grid $\bm{\mathcal{E}}$, following NeRFShop [16].
To facilitate region selection, the user scribbles on the projection of the 3D scene on the screen. Subsequently, for each selected pixel, a ray is cast and the estimated ray termination ${\bm{x}}_{\text{term}}$ is computed with Eq. 5. Next, we map each ${\bm{x}}_{\text{term}}$ to the nearest voxel in our edit grid $\bm{\mathcal{E}}$ and set the corresponding bit in the underlying bitfield. For intuitive selection, we support user-controlled region growing from the selected voxels, by adding the neighboring voxels to a growing queue ${\mathcal{G}}$. During region growing, we add the current voxel to $\bm{\mathcal{E}}$ and its neighbors to ${\mathcal{G}}$ if the $\bm{\mathcal{O}}$ is set. Through this workflow, we offer an intuitive method for content selection within iNGP’s hybrid representation.
Figure 2: Overview of LAENeRF: Given the estimated ray terminations ${\bm{x}}_{\text{term}}$ for a region specified by $\bm{\mathcal{E}}$, we learn a mapping from positions to weights ${\hat{{\bm{w}}}}$ and offsets ${\hat{\bm{\delta}}}$. We compose the color ${\hat{{\bm{c}}}}$ using these latent outputs and a learnable color palette ${\hat{\bm{\mathcal{P}}}}$ and supervise with a content loss ${\mathcal{L}_{\text{content}}}$ and optional style losses ${\mathcal{L}_{\text{style}}},{\mathcal{L}_{\text{TV}}},{\mathcal{L}_{\text{depth-disc}}}$ to obtain a unified approach which supports recoloring and stylization.
@@ 4 LAENeRF
Our key insight is that we can reduce computational requirements by optimizing a lightweight, NeRF-like network given 3D positions ${\bm{x}}_{\text{term}}$, which represents a local, edited region. Starting from $\bm{\mathcal{E}}$ defined by the user, we extract ${\bm{x}}_{\text{term}}$ for all training views, optimize LAENeRF and transfer the changes to $\Theta_{\text{NeRF}}$ with an efficient distillation step. We present our network architecture in Fig. 2. ${\bm{x}}_{\text{term}}$ is featurized using a trainable multi-resolution hash grid [29]. The encoded input $\xi({\bm{x}}_{\text{term}})$ is subsequently passed to shallow MLPs to predict per-ray barycentric weights ${\hat{{\bm{w}}}}\in[0,1]^{{N_{{\hat{\bm{\mathcal{P}}}}}}}$ and view-dependent offsets ${\hat{\bm{\delta}}}\in[-1,1]^{3}$, where ${N_{{\hat{\bm{\mathcal{P}}}}}}$ denotes the size of the learnable color palette ${\hat{\bm{\mathcal{P}}}}\in\mathbb{R}^{{N_{{\hat{\bm{\mathcal{P}}}}}}\times 3}$. Finally, we compose our output color as
@@
We featurize ${\bm{d}}$ using a spherical harmonics encoding [8, 44] and use this as an additional input to our offset network.
@@
As can be seen in Fig. 2LAENeRF learns a mapping from ${\bm{x}}_{\text{term}}$ to estimated output color ${\hat{{\bm{c}}}}$. However, we only require ${\bm{x}}_{\text{term}}$ of rays ${\bm{r}}$ which intersect $\bm{\mathcal{E}}$. To correctly handle occlusions and mitigate errors due to the edit grid resolution, we require the alpha accumulated inside the edit grid $T_{\alpha}$ to be larger than ${\tau_{\text{edit}}}=0.5$, where we define $T_{\alpha}$ as
[…]
with $\mathbbm{1}(\cdot)$ denoting the indicator function. To obtain accurate depth estimates $0pt$, requiring the full accumulation of alpha, we simultaneously raymarch through both $\bm{\mathcal{E}}$ and the occupancy grid $\bm{\mathcal{O}}$, computing $0pt$ within the occupancy grid.
@@
As is common in image style transfer [9, 17], we use separate losses for content and style:
${\mathcal{L}_{\text{content}}}$ $=\|{\hat{{\bm{c}}}}-\mathcal{C}({\bm{r}})\|_{2}^{2},$ (8)
@@
where ${{\bm{s}}}$ denotes an arbitrary style image and ${\bm{G}}(\cdot)$ denotes the gram matrix [9] of a feature extracted from a semantic encoder, e.g. VGG19 [36]. Crucially, we can perform photorealistic recoloring by setting ${\lambda}_{\text{style}}=0$, where LAENeRF learns to reconstruct the selected region due to Eq. 8.
@@
We notice that when we perform non-photorealistic stylization of 3D regions using only ${\mathcal{L}_{\text{content}}}$ and ${\mathcal{L}_{\text{style}}}$, small structures are often eliminated in favor of a consistent stylization. In addition, as imperfect geometry reconstruction from our pre-trained $\Theta_{\text{NeRF}}$ leads to noise in ${\bm{x}}_{\text{term}}$, we need additional regularization to encourage smooth, low-noise outputs. To facilitate detailed, geometry-aware stylization, we introduce two novel losses, which condition LAENeRF on the estimated geometry of $\Theta_{\text{NeRF}}$.
First, we encourage LAENeRF to limit noise in regions without depth discontinuities using a depth guidance image:
@@
where $\nabla{\hat{{\bm{c}}}}$ denotes the image gradients of ${\hat{{\bm{c}}}}$ in $x/y$-direction, i.e. $(\nabla{\hat{{\bm{c}}}})_{i,j}=\|{\hat{{\bm{c}}}}_{i+1,j}-{\hat{{\bm{c}}}}_{i,j}\|_{2}^{2}+\|{\hat{{\bm{c}}}}_{i,j+1}-{\hat{{\bm{c}}}}_{i,j}\|_{2}^{2}$. This loss term remedies noisy prediction for ${\bm{x}}_{\text{term}}$, but does not preserve fine, geometric structures sufficiently. Hence, we introduce another loss, which is minimized when image gradients are placed in regions of depth discontinuities:
@@
To prevent extreme solutions with a high-frequency offset function, we use the offset loss from Aksoy et al. [1]:
@@
To transfer the local changes encoded in LAENeRF to the pre-trained radiance field, we require an additional fine-tuning step. First, we obtain a modified training dataset, where we compose the target color as ${T_{\alpha}}{\hat{{\bm{c}}}}+(1-{T_{\alpha}})\mathcal{C}({\bm{r}})$. Thus, rays which did not intersect $\bm{\mathcal{E}}$ use the ground truth color $\mathcal{C}({\bm{r}})$ during distillation, which effectively mitigates background artefacts.
@@
where ${\tau_{\text{dist}}}$ is a hyperparameter controlling the size of the transition, which we usually set to $1\times 10^{-2}$, depending on the scale of the scene. As can be seen in Fig. 3, we interpolate the original color palette ${\hat{\bm{\mathcal{P}}}}$ and the user-modified color palette ${\hat{\bm{\mathcal{P}}}}_{\text{mod}}$ based on ${d_{\text{trans}}}$ to get a realistic color transition. For stylization, we additionally introduce a separate loss, which adds to the content loss when ${d_{\text{trans}}}\in(0,1]$:
@@
We build our approach on torch-ngp [40] and use a resolution of $128^{3}$ for our edit grid $\bm{\mathcal{E}}$, following iNGP [29]. We train LAENeRF for $1\times 10^{5}$ iterations with previews available after $\sim 20$s, and then distill to $\Theta_{\text{NeRF}}$ for $7\times 10^{4}$ iterations, which takes no longer than 5 minutes in total on an NVIDIA RTX 4090. For LAENeRF, we use a learning rate of $1\times 10^{-3}$ for all components, except for ${\hat{\bm{\mathcal{P}}}}$, where we use $1\times 10^{-2}$. We initialize ${N_{{\hat{\bm{\mathcal{P}}}}}}=8$, as the use of ${\mathcal{L}_{\text{smooth}}}$ requires a large number of base colors. $1.5\times 10^{3}$ iterations before training is finished, we remove color palettes which do not contribute significantly. For our style loss, we use a VGG19 backbone [36] and utilize features from conv5, conv6, conv7 for computing ${\mathcal{L}_{\text{style}}}$. When performing stylization, we use ${\lambda}_{\text{style}}=1.3\times 10^{2},{\lambda}_{\text{TV}}=1\times 10^{-4},{\lambda}_{\text{depth-disc}}=5\times 10^{-4},{\lambda}_{\text{weight}}=1\times 10^{-7},{\lambda_{\text{offset}}}=5\times 10^{-5},{\lambda}_{\text{smooth}}=1\times 10^{-3}$. If we want to perform photorealistic recoloring, we set ${\lambda}_{\text{style}}={\lambda}_{\text{TV}}={\lambda}_{\text{depth-disc}}=0$.